Wildhorse Company has decided to introduce a new product. The new product can be manufactured by either a capital-intensive mDetermine the annual unit sales volume at which Wildhorse Company would be indifferent between the two manufacturing methods.
Answer #1
Capital intensive manufacturing method
breakeven point = fixed cost/contribution margin p.u
contribution margin
sale price 31
less: variable cost 16
(5+6+3+2)
contribution margin 15
Fixed cost = (2278000+482000) = 2760000
Breakeven point = 2760000/15 = 184000
labor intensive manufacturing method
fixed cost = (1410000+482000) = 1892000
contribution margin = 31-(5.50+8+4.50+2) = 11
Breakeven point = 1892000/11 = 172000 units
2) assume volume units sold = x
profit= contribution margin-fixed cost
capital intensive manufacturing method= 15x-2760000
labor intensive manufacturing method = 11x-1892000
15x-2760000 = 11x-1892000
4x = 868000
x = 217000 Units